# Improving Speech Emotion Recognition with Mutual Information Regularized Generative Model

### v1

@@ 3.2.2 Information Maximizing Generative Adversarial Network (InfoGAN)
Note that the same weights from the baseline model is used to enforce latent variable to be coded with emotional information. If not, InfoGAN training objective will code arbitrary information of variations within the dataset (similar to clustering results). Similar approach is taken for latent variable $t$, but use InfoNCE loss instead of mean squared error loss, as we are optimizing embedding model rather than regression model.
$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t}$ (17)
$\mathcal{L}_{I_{t}}$ $=-log\frac{\exp\left(sim(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(sim(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (18)
Again, that the same weights from the baseline model is used to enforce latent variable to be coded with textual information. Furthermore, using the weights used in baseline model allows us to maximize mutual information not only with generated samples but also with real samples.
3.3 Data augmentation for SER
After InfoGAN is successfully trained, we can train emotion prediction model by training data samples, together with generated data samples that resembles their distribution. As final stage, all modules, except emotion classification module, are frozen.
The case for SER is rather straightforward. After extracting $h$ from training dataset, we generate same number of $\hat{h}$, and then fed them through same emotion classification layer and compute loss to further fine-tune them. In effect, the emotion predictor has been trained with dataset that has size of double from original training dataset.
For the case of multimodal SER, $h$, $\hat{h}$, $t$ and $t^{\prime}$ has to be prepared. And new classification layer that takes concatenated feature of audio feature and text feature has to be defined. Then we can compose inputs with all possible concatentation with original and generated features. Thus, we can achieve the data sample size increase by factor of four.
@@ 4.1.3 MSP-Podcast [28]
4.2 Data augmentation for SER
4.2.1 Baseline model
@@
4.3 InfoGAN architecture
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly we adopted mix-up strategy for training generator and discriminator as proposed in [6], for stable training of generator.
4.4 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [30] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [29], that is ’roberta-uncased-large’ version. Fof multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 3.0 instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same.
@@ 5.1 SER experiment
We experimented our method on commonly used dataset IEMOCAP. The experiment followed leave-one-speaker-out cross validation scheme. Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table VI as follows.
@@ 5.2.3 MSP-Podcast
Audio Text Audio + Text
emoDARTS [36] 61.15 - -
Ours 56.76 49.01 60.61
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [36], which uses neural architecture search and only audio input. Our method performed up to similar level of success to emoDARTS. Compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [12], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, out data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
@@ 6 Conclusion
We proposed a new framework of data augmentation for speech emotion recognition (SER) using GAN. Unlike previous approaches, we employ the combination of mutual information regularization and cross-modality transfer (from text to be specific). The advantage of mutual information regularization is: mutual information regularizing loss can serve as an observable metric to depict the quality of generated data sample, in the form of dependency measure. Furthermore, we propose multimodal data augmentation using our framework. Previous works were limited in using cross-modal information transfer or cross-modal generative model for supplementing noisy of missing modality inputs. In our work, we propose to use our cross-modal generative model to generate more multimodal inputs by preparing combination of real inputs and generated inputs. We conducted experiments to test our framework in both, SER and multimodal SER cases. The framework was implemented to generate audio features or text features that is fed into final classification layer. Our method was evaluated on IEMOCAP, MSP-IMPROV and MSP-Podcast. Thanks to our method, we observed decent amount of improvement. We conclude, that when one has two separate decent models to predict emotion from audio and text, our framework can be implemented to generate more inputs further improve the performance in multimodal emotion recognition. Additionally, we discovered that with mutual information reglarization, one can generate more decent inputs without any information from other modality.

### v2

@@ 3.2 Mutual Information Regularised Feature Generation
[…]
We re-use the pre-trained prediction layers from the baseline model ($W_{y},b_{y}$ and $W_{t},b_{t}$), freezing them during this stage. This weight-sharing strategy ensures the latent codes $c$ and $t$ are forced to represent emotional and textual information, respectively.
The mutual information losses are defined as:
Emotion Prediction Loss: A cross-entropy loss ensuring the generated feature $\hat{h}$ predicts the correct emotion label $c$:
$\hat{y}_{g}$ $=\text{softmax}(W_{y}\hat{h}+b_{y}),$ (10)
[…]
Text Prediction Loss: An InfoNCE loss that attracts the projected text embedding from $\hat{h}$ to the true conditioning text embedding $t$:
$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t},$ (12)
$\mathcal{L}_{I_{t}}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (13)
The full objective for the InfoGAN is:
$\min_{G}\max_{D}V(D,G)-\lambda_{y}\mathcal{L}_{I_{y}}-\lambda_{t}\mathcal{L}_{I_{t}},$ (14)
where $\lambda_{y}$ and $\lambda_{t}$ are hyperparameters controlling the strength of the mutual information regularisation.
3.3 Data Augmentation and Classifier Fine-tuning
In the final stage, we leverage the trained generator for data augmentation. All feature extraction and generation modules are frozen, and only the classification layers are fine-tuned.
Unimodal SER: We generate a set of synthetic audio features $^{\hat{h}}$ equal in size to the original training set. The classifier is then fine-tuned on the combined set $^{h}\cup{\hat{h}}$, effectively doubling the training data.
[…]
This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent samples that directly address the core challenges of data scarcity in SER.
@@ 4.1.3 MSP-Podcast [35]
4.2 Model configuration
4.2.1 Data augmentation for SER
@@
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [4], for stable training of generator.
4.2.2 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [36] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [32], that is ’roberta-uncased-large’ version. For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 2.0 [37] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same. Multimodal SER classification layer takes concatenated feature of both modalities. The classifier consists linear layer, one hidden layer (256) or two hidden layers (256-128), for IEMOCAP, MSP-IMPROV or MSP-Podcast respectively.
4.2.3 Hyperparameters
We used Adam optimizer with 1e41e-4 learning rate for training. We employed different weights per class during cross entropy loss computation due to class imbalance, for MSP-IMPROV and MSP-Podcast. The weights were inverse proportional to the total numbers of samples per each class. Lora [38] was adopted when Llama was trained for text emotion recognition. The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
@@ 5.1 SER experiment
We experimented our method on commonly used dataset IEMOCAP. The experiment followed leave-one-speaker-out cross validation scheme. Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table II as follows.
@@ 5.2.3 MSP-Podcast
emoDARTS [45] 61.15 ± 2.41
Ours 60.00 ± 0.72
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [45], which uses neural architecture search and only audio input. Our method performed better than emoDARTS in terms of variance. Furthermore, compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [29], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, our data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
5.3 Qualitative analysis of generated features
Figure 2: T-SNE plot for both [real and generated combined] (left), generated (center) and real (right) audio features
We visualized both generated features from InfoGAN and real audio features for qualitative analysis. The data samples from test dataset of MSP-Podcast was used and same number of generated samples were prepared. Scatter plots of the samples, after transformation via t-sne [46], were produced and analysed. It is clearly observable that samples from different classes are placed on different separable regions (refer to Fig.2).
@@ 6 Conclusion
We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation. Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated samples and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
The key insights from our work are threefold. First, mutual information regularisation provides both a training objective and a measurable quality metric, enabling controllable generation of emotionally consistent features. Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression. Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text samples that significantly enhance multimodal SER performance.
Extensive experiments across three diverse datasets validate our approach, demonstrating consistent improvements over existing methods in both unimodal and multimodal settings. The ablation studies further confirm the individual contributions of cross-modal alignment and mutual information regularization to the overall framework’s effectiveness.
Despite these promising results, the current work has limitations. The framework assumes access to aligned speech–text pairs (for obtaining embeddings) and relies on a pretrained SER encoder, which may not be available for all languages or domains. Moreover, the GAN synthesises features rather than raw waveforms, so the realism of generated speech may be limited. Future work will address these issues by exploring end-to-end emotional speech synthesis (e.g. using text-to-speech or diffusion-based generative models) to produce waveform-level samples, potentially improving audio fidelity. We also plan to investigate unsupervised or weakly supervised extensions that reduce reliance on transcripts, and to apply the method to multilingual or continuous-emotion settings. These directions may further advance data-efficient SER by leveraging generative models to capture the nuance of emotional speech.